OpenStudy (anonymous):

Use DeMoivre's Theorem to find all complex solutions. Express in rectangular form. Help??? x^3-i=0

OpenStudy (anonymous):

Will give a medal for any help!

OpenStudy (anonymous):

plz just need some advice...

OpenStudy (anonymous):

try this one: http://college.cengage.com/mathematics/larson/elementary_linear/4e/shared/downloads/c08s3.pdf I'm looking in to it... its been a while :)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Express the result in rectangular form. ... is equivalent to the polar form

OpenStudy (anonymous):

ok I got that...

OpenStudy (anonymous):

this one is good! - check example #4: http://mathonweb.com/help_ebook/html/complex_2.htm

OpenStudy (anonymous):

Let\[z ^{3}=i\]\[z _{1}=\cos(\pi/4)+i \sin (\pi/4)\]\[=(\sqrt{2}/2)+(\sqrt{2}/2)i\]

OpenStudy (anonymous):

\[z _{2}=\cos (5\pi/4)+i \sin (5\pi/4)\]\[=(-\sqrt{2}/2)-(\sqrt{2}/2)i\]

OpenStudy (anonymous):

\[z _{3}=\cos (9\pi/4)+i \sin (9\pi/4)\]=same as z(1)